# REOBench: Benchmarking Robustness of Earth Observation Foundation Models

### v1

@@ Abstract
Earth observation foundation models have shown strong generalization across multiple Earth observation tasks, but their robustness under real-world perturbations remains underexplored. To bridge this gap, we introduce REOBench, the first comprehensive benchmark for evaluating the robustness of Earth observation foundation models across six tasks and twelve types of image corruptions, including both appearance-based and geometric perturbations. To ensure realistic and fine-grained evaluation, our benchmark focuses on high-resolution optical remote sensing images, which are widely used in critical applications such as urban planning and disaster response. We conduct a systematic evaluation of a broad range of models trained using masked image modeling, contrastive learning, and vision-language pre-training paradigms. Our results reveal that ❶ existing Earth observation foundation models experience significant performance degradation when exposed to input corruptions. ❷ The severity of degradation varies across tasks, model architectures, backbone sizes, and types of corruption, with performance drop varying from less than 1% to over 20%. ❸ Vision-language models show enhanced robustness, particularly in multimodal tasks. REOBench underscores the vulnerability of current Earth observation foundation models to real-world corruptions and provides actionable insights for developing more robust and reliable models. Code and data are publicly available at https://github.com/lx709/REOBench.
@@ 2.2 Definition of Corruption Robustness
We formalize the corruption robustness of EOFMs as its ability to maintain task performance in the presence of realistic geophysical and sensor-induced degradations frequently encountered in Earth observation. Let $f:\mathcal{X}\to\mathcal{Y}$ denote an EOFMs that maps an input image $x\in\mathcal{X}$ to a label $y\in\mathcal{Y}$, with $(x,y)$ sampled from an underlying geospatial data-generating distribution $\mathcal{D}$. We define a set of corruptions $\mathcal{C}=\{c_{1},\dots,c_{K}\}$, where each $c_{k}\in\mathcal{C}$ represents a physically plausible corruption operator, such as haze, cloud occlusion, or sensor noise. Each corruption occurs with a non-zero prevalence $\mathbb{P}_{\mathcal{C}}(c_{k})>0$. To quantify robustness, we introduce the task performance drop $\Delta_{\text{TP}}$, which captures the expected degradation in performance when the model is evaluated on corrupted inputs:
$\Delta_{\text{TP}}\;=\;\mathbb{P}_{(x,y)\sim\mathcal{D}}\!\left[f(x)=y\right]\;-\;\mathbb{E}_{c\sim\mathcal{C}}\!\left[\mathbb{P}_{(x,y)\sim\mathcal{D}}\!\left[f(c(x))=y\right]\right].$ (1)
A smaller $\Delta_{\text{TP}}$ value indicates greater robustness, as it reflects a smaller degradation in performance when the model encounters realistic corruptions.
@@ 3.1 Implementation Details
For the scene classification task, we take the backbone from all pretrained foundation models and append a single linear layer after the backbone for classification. For the semantic segmentation (resp. detection) task, we follow RVSA [46] to use UpperNet [55] (resp. Oriented R-CNN [56]) as the segmentor (resp. detector) and replace its backbone with that from pretrained foundation models. Following RVSA [46], we take the feature maps from block 4, block 6, block 8, and block 12 to form a feature pyramid through upsampling and downsampling. We then train all comparing models for 12 epochs with an initial learning rate of 1e-5. We decay the learning rate by a factor of 0.1 at epochs 8 and 11.
@@ 3.2 Scene Classification
From Table 1, we can draw the following findings: 1) All benchmark methods suffer from serious performance under image corruptions for the scene classification tasks, especially for MIM-based methods. 2) CL- and LLM-based methods are more robust towards image corruptions than MIM-based methods. This is probably because CL- and LLM-based methods are trained by matching image-text pairs in a shared embedding space, learning high-level semantic features less sensitive to low-level corruptions. In contrast, MIM-based methods are trained by reconstructing pixel- or token-level details, making them sensitive to local corruptions. Specifically, VHM [52] achieves the least performance drop under image corruptions. 3) CL-based methods perform better than MIM and LLM-based methods on the scene classification task, for both clean and noisy images. Specifically, GeoRSCLIP [10] achieves the best scene classification performance under image corruptions.
Table 1: Scene classification performance on AID dataset across different image perturbations. ${zs}$ denotes zero-shot evaluation.
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Rotate Salt Scale Translate Avg $\Delta_{\text{TP}}$
@@
SATLAS [57] Swin-B 90.85 82.54 84.32 73.36 67.23 78.10 79.16 80.46 32.44 72.54 77.56 72.54 88.54 74.07 16.78
SatMAE [5] ViT-L 72.05 44.82 59.58 67.26 46.49 71.33 71.25 28.31 63.85 69.15 70.45 59.74 66.12 59.86 12.19
Scale-MAE [7] ViT-L 75.75 51.80 72.65 39.60 43.69 31.65 46.31 55.24 17.49 66.15 47.27 61.58 69.84 50.27 25.48
RVSA [46] ViT-B 84.60 56.84 77.33 56.07 53.14 53.53 32.51 49.19 23.45 76.88 35.12 71.78 77.22 55.26 29.35
SatMAE++ [47] ViT-L 91.35 64.62 82.64 62.69 60.70 48.23 76.98 62.56 29.43 85.49 73.22 75.79 87.61 67.50 23.85
@@
RemoteCLIPzs [9] ViT-L 81.10 78.32 80.64 73.91 79.43 76.83 76.72 80.10 57.02 82.80 70.90 68.39 80.72 75.48 5.62
RemoteCLIP [9] ViT-B 96.85 90.80 95.36 91.13 88.96 89.18 94.25 87.46 63.75 96.22 91.43 83.62 95.42 88.97 7.89
RemoteCLIP [9] ViT-L 95.45 93.11 93.80 88.77 94.21 92.47 94.20 93.37 74.45 95.01 86.99 83.37 94.06 90.32 5.13
GeoRSCLIPzs [10] ViT-L 66.05 62.41 65.47 60.45 64.41 62.03 62.32 62.24 44.20 65.52 58.88 52.59 64.25 60.40 5.65
GeoRSCLIP [10] ViT-B 96.90 93.59 96.04 91.01 93.34 92.60 92.99 92.91 57.78 95.70 88.22 75.70 93.87 88.65 8.25
GeoRSCLIP [10] ViT-L 97.40 96.27 96.45 92.28 95.62 96.09 95.68 96.00 71.03 97.20 92.75 77.16 95.15 91.80 5.59
@@
GeoChat [12] ViT-L 65.85 64.67 65.26 60.71 64.61 63.32 62.34 64.54 48.68 65.05 62.32 56.21 62.91 61.72 4.13
LHRS-Bot [50] ViT-L 87.75 87.38 86.82 78.53 85.95 84.94 82.62 87.62 67.76 87.31 76.73 79.07 86.71 82.79 4.96
RS-LLaVA [51] ViT-L 67.55 65.36 68.95 63.05 67.69 65.73 63.13 65.97 43.86 66.55 68.24 54.89 63.40 63.07 4.48
SkySenseGPT [53] ViT-L 87.35 87.72 87.91 79.66 87.67 84.78 83.08 87.71 63.11 86.86 83.61 75.49 85.25 82.74 4.61
VHM [52] ViT-L 80.60 79.81 80.67 76.10 80.82 78.81 76.92 79.55 59.74 80.47 75.43 72.57 79.73 76.72 3.88
@@ 3.3 Semantic Segmentation
The ISPRS Potsdam dataset provides both non-eroded and eroded labels, corresponding to annotations with and without object boundaries, respectively. In our experiments, we use the non-eroded labels for model evaluation and report the mean IoU for MIM- and CL-based methods. We omit LLM-based methods for semantic segmentation due to the lack of open-source MLLMs for this task. From Table 2, we can draw the following findings: 1) both MIM- and CL-based methods suffer from serious performance under image corruptions for the object detection task, with a mIoU drop of more than 5%. 2) MIM-based methods achieve much better performance than CL-based methods on clean and noisy images for semantic segmentation in remote sensing images. This is probably because MIM-based methods can capture local details by pixel reconstruction, while CL-based methods align global visual embeddings with text features, thus losing fine-grained details. Specifically, ScaleMAE [7] achieves the best segmentation performance under image corruption, with an average mIoU of 60.02%. 3) MIM-based methods suffer from a more serious performance drop under image corruptions than CL-based methods. Specifically, GeoRSCLIP [10] achieves the best robustness across image corruptions, with a drop of 5.15% of mIoU under corruptions.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Rotate Salt Scale Translate Avg $\Delta_{\text{TP}}$
@@
SatMAE [5] ViT-L 59.51 50.18 37.39 48.23 51.15 57.89 41.83 44.95 57.52 56.02 36.07 54.81 59.09 49.59 9.92
ScaleMAE [7] ViT-L 68.92 64.37 65.43 49.96 41.84 64.86 54.89 63.89 64.49 64.51 52.12 65.45 68.38 60.02 8.90
RVSA [46] ViT-B 69.82 64.71 65.67 47.99 45.34 66.89 48.89 61.52 65.25 65.58 45.26 66.87 69.23 59.43 10.39
SatMAE++ [47] ViT-L 62.68 53.91 59.06 49.74 53.94 60.38 44.32 48.80 60.44 58.34 39.45 58.13 61.94 54.04 8.64
@@
RemoteCLIP [9] ViT-B 50.28 42.32 45.27 39.33 37.78 50.26 48.46 36.61 50.06 46.48 46.91 48.39 49.63 45.12 5.16
RemoteCLIP [9] ViT-L 56.69 54.51 52.73 43.24 51.19 50.82 45.12 50.47 51.82 53.68 38.98 54.59 56.53 50.31 6.38
GeoRSCLIP [10] ViT-B 51.44 42.89 46.41 40.37 38.64 51.35 49.79 38.56 51.15 47.89 48.24 49.28 50.87 46.29 5.15
GeoRSCLIP [10] ViT-L 56.81 54.97 52.53 43.28 41.37 50.49 42.7 49.41 51.36 53.54 36.98 54.66 56.64 48.99 7.82
@@ 3.4 Object Detection
We evaluate robustness on the corrupted images from the DIOR dataset. We report mAP for MIM- and CL-based methods. We omit LLM-based methods for the object detection task due to the lack of open-source LLMs for this task. From Table 3, we can draw the following findings: 1) both MIM- and CL-based methods suffer from serious performance under image corruptions, with a mAP drop of more than 5%. 2) MIM- and CL-based methods achieve comparable performance for object detection on clean and noisy images. RVSA [46] attains the highest mAP of 70.96% on clean images but experiences a substantial performance decline under image corruptions. 3) MIM-based and CL-based methods exhibit a similar degree of performance degradation when subjected to image corruptions. Among them, GeoRSCLIP [10] demonstrates the most robust detection performance under noisy conditions.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Rotate Salt Scale Translate Avg $\Delta_{\text{TP}}$
@@
SatMAE [5] ViT-L 62.30 56.84 57.86 55.80 58.36 55.38 58.44 59.34 56.92 56.60 53.76 51.58 60.90 56.82 5.49
ScaleMAE [7] ViT-L 70.20 64.80 65.98 62.50 64.46 62.58 63.82 66.10 63.08 63.44 60.50 53.08 68.26 63.22 6.98
RVSA [46] ViT-B 70.96 60.59 65.02 61.58 64.60 62.35 62.87 63.98 62.88 64.04 56.61 55.97 69.69 62.51 8.45
SatMAE++ [47] ViT-L 65.20 59.44 61.02 60.30 59.88 59.66 61.06 61.72 59.56 59.14 58.64 48.48 64.70 59.47 5.73
@@
RemoteCLIP [9] ViT-B 60.40 56.72 56.28 56.78 54.56 53.68 57.36 55.90 53.42 54.54 54.40 44.92 59.72 54.86 5.54
RemoteCLIP [9] ViT-L 70.20 66.52 66.62 63.84 65.40 63.62 63.68 66.76 62.66 63.52 59.16 57.42 68.64 63.99 6.21
GeoRSCLIP [10] ViT-B 60.20 56.28 56.04 56.08 55.46 53.38 56.92 55.50 53.38 53.98 53.48 46.98 59.32 54.73 5.47
GeoRSCLIP [10] ViT-L 69.80 66.12 65.34 65.34 64.96 63.62 62.90 66.04 62.02 62.68 56.04 57.40 68.10 63.38 6.42
@@ 3.5 Image Captioning
As shown in the upper part of Table 4, all models experience performance degradation under noisy conditions. Among them, the Falcon [54] model achieves the best overall performance, significantly outperforming other methods on both clean and corrupted images, with an average performance drop of 3.89%. In contrast, RS-LLaVA [51] demonstrates the strongest robustness to image corruptions, exhibiting the smallest decrease in CLAIR score, with only a 1.04% drop. Additionally, we present results for GeoChat [12], fine-tuned on the VRSBench training set, as shown in the lower part of Table 4. The fine-tuned GeoChat model on the target dataset exhibits significantly improved performance compared to its zero-shot counterpart, but suffers from a larger performance drop under image perturbations.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Salt Avg $\Delta_{\text{TP}}$
@@
GeoChat [12] ViT-L 41.39 40.06 40.45 37.65 40.20 39.76 38.48 40.38 39.92 37.61 39.59 1.80
SkySenseGPT [53] ViT-L 48.29 47.21 46.64 44.22 46.25 45.52 44.97 46.14 45.13 44.36 45.60 2.69
VHM [52] ViT-L 52.02 50.19 50.82 50.26 50.57 51.22 50.46 50.39 50.72 49.48 50.46 1.56
RS-LLaVA [51] ViT-L 51.30 51.15 50.43 51.78 50.54 52.01 47.84 50.57 49.88 48.12 50.26 1.04
Falcon [54] DaViT-B 61.90 59.98 60.09 57.13 59.48 57.43 56.31 59.85 59.94 51.83 58.01 3.89
GeoChatft [12] ViT-L 71.26 69.00 68.93 66.60 69.45 68.63 67.83 69.98 69.02 63.87 68.15 3.11

### v2

@@ Abstract
Earth observation foundation models have shown strong generalization across multiple Earth observation tasks, but their robustness under real-world perturbations remains underexplored. To bridge this gap, we introduce REOBench, the first comprehensive benchmark for evaluating the robustness of Earth observation foundation models across six tasks and twelve types of image corruptions, including both appearance-based and geometric perturbations. To ensure realistic and fine-grained evaluation, our benchmark focuses on high-resolution optical remote sensing images, which are widely used in critical applications such as urban planning and disaster response. We conduct a systematic evaluation of a broad range of models trained using masked image modeling, contrastive learning, and vision-language pre-training paradigms. Our results reveal that ❶ existing Earth observation foundation models experience significant performance degradation when exposed to input corruptions. ❷ The severity of degradation varies across tasks, model architectures, backbone sizes, and types of corruption, with performance drop varying from less than 1% to over 25%. ❸ Vision-language models show enhanced robustness, particularly in multimodal tasks. REOBench underscores the vulnerability of current Earth observation foundation models to real-world corruptions and provides actionable insights for developing more robust and reliable models. Code and data are publicly available at https://github.com/lx709/REOBench.
@@ 2.2 Definition of Corruption Robustness
We formalize the corruption robustness of EOFMs as its ability to maintain task performance in the presence of realistic geophysical and sensor-induced degradations frequently encountered in Earth observation. Let $f:\mathcal{X}\to\mathcal{Y}$ denote an EOFMs that maps an input image $x\in\mathcal{X}$ to a label $y\in\mathcal{Y}$, with $(x,y)$ sampled from an underlying geospatial data-generating distribution $\mathcal{D}$. We define a set of corruptions $\mathcal{C}=\{c_{1},\dots,c_{K}\}$, where each $c_{k}\in\mathcal{C}$ represents a physically plausible corruption operator, such as haze, cloud occlusion, or sensor noise. Each corruption occurs with a non-zero prevalence $\mathbb{P}_{\mathcal{C}}(c_{k})>0$. To quantify robustness, we define the relative task performance drop ($\mathcal{R}_{\text{TP}}$), which measures the degradation in model performance under corrupted inputs relative to its accuracy on clean data:
$\mathcal{R}_{\text{TP}}\;=\;\frac{\mathbb{P}_{(x,y)\sim\mathcal{D}}\!\left[f(x)=y\right]-\mathbb{E}_{c\sim\mathcal{C}}\!\left[\mathbb{P}_{(x,y)\sim\mathcal{D}}\!\left[f(c(x))=y\right]\right]}{\mathbb{P}_{(x,y)\sim\mathcal{D}}\!\left[f(x)=y\right]}.$ (1)
A smaller $\mathcal{R}_{\text{TP}}$ indicates greater robustness, as it reflects less relative degradation when encountering corrupted data.
@@ 3.1 Implementation Details
For the scene classification task, we take the backbone from all pretrained foundation models and append a single linear layer after the backbone for classification. For the semantic segmentation (resp. detection) task, we follow RVSA [46] to use UpperNet [55] (resp. Oriented R-CNN [56]) as the segmentor (resp. detector) and replace its backbone with that from pretrained foundation models. Following RVSA [46], we build a feature pyramid from blocks 4, 6, 8, and 12 using up/down-sampling. All models are trained for 12 epochs with an initial learning rate of 1e-5, decayed by 0.1 at epochs 8 and 11.
@@ 3.2 Scene Classification
From Table 1, we can draw the following findings: 1) All benchmark methods suffer from serious performance under image corruptions for the scene classification tasks, especially for MIM-based methods. 2) CL- and LLM-based methods are more robust towards image corruptions than MIM-based methods. This is probably because CL- and LLM-based methods are trained by matching image-text pairs in a shared embedding space, learning high-level semantic features less sensitive to low-level corruptions. In contrast, MIM-based methods are trained by reconstructing pixel- or token-level details, making them sensitive to local corruptions. Specifically, VHM [52] achieves the least performance drop under image corruptions. 3) CL-based methods usually perform better than MIM and LLM-based methods on the scene classification task, for both clean and noisy images. Specifically, GeoRSCLIP [10] achieves the best scene classification performance under image corruptions.
Table 1: Scene classification performance on AID dataset across different image perturbations. zs{zs} denotes zero-shot evaluation.
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Rotate Salt Scale Translate Avg $\mathcal{R}_{\text{TP}}$
@@
SATLAS [57] Swin-B 90.85 82.54 84.32 73.36 67.23 78.10 79.16 80.46 32.44 72.54 77.56 72.54 88.54 74.07 18.47
SatMAE [5] ViT-L 72.05 44.82 59.58 67.26 46.49 71.33 71.25 28.31 63.85 69.15 70.45 59.74 66.12 59.86 16.92
Scale-MAE [7] ViT-L 75.75 51.80 72.65 39.60 43.69 31.65 46.31 55.24 17.49 66.15 47.27 61.58 69.84 50.27 33.64
RVSA [46] ViT-B 84.60 56.84 77.33 56.07 53.14 53.53 32.51 49.19 23.45 76.88 35.12 71.78 77.22 55.26 34.69
SatMAE++ [47] ViT-L 91.35 64.62 82.64 62.69 60.70 48.23 76.98 62.56 29.43 85.49 73.22 75.79 87.61 67.50 26.11
@@
RemoteCLIPzs [9] ViT-L 81.10 78.32 80.64 73.91 79.43 76.83 76.72 80.10 57.02 82.80 70.90 68.39 80.72 75.48 6.93
RemoteCLIP [9] ViT-B 96.85 90.80 95.36 91.13 88.96 89.18 94.25 87.46 63.75 96.22 91.43 83.62 95.42 88.97 8.15
RemoteCLIP [9] ViT-L 95.45 93.11 93.80 88.77 94.21 92.47 94.20 93.37 74.45 95.01 86.99 83.37 94.06 90.32 5.38
GeoRSCLIPzs [10] ViT-L 66.05 62.41 65.47 60.45 64.41 62.03 62.32 62.24 44.20 65.52 58.88 52.59 64.25 60.40 8.55
GeoRSCLIP [10] ViT-B 96.90 93.59 96.04 91.01 93.34 92.60 92.99 92.91 57.78 95.70 88.22 75.70 93.87 88.65 8.51
GeoRSCLIP [10] ViT-L 97.40 96.27 96.45 92.28 95.62 96.09 95.68 96.00 71.03 97.20 92.75 77.16 95.15 91.80 5.74
@@
GeoChat [12] ViT-L 65.85 64.67 65.26 60.71 64.61 63.32 62.34 64.54 48.68 65.05 62.32 56.21 62.91 61.72 6.27
LHRS-Bot [50] ViT-L 87.75 87.38 86.82 78.53 85.95 84.94 82.62 87.62 67.76 87.31 76.73 79.07 86.71 82.79 5.65
RS-LLaVA [51] ViT-L 67.55 65.36 68.95 63.05 67.69 65.73 63.13 65.97 43.86 66.55 68.24 54.89 63.40 63.07 6.63
SkySenseGPT [53] ViT-L 87.35 87.72 87.91 79.66 87.67 84.78 83.08 87.71 63.11 86.86 83.61 75.49 85.25 82.74 5.28
VHM [52] ViT-L 80.60 79.81 80.67 76.10 80.82 78.81 76.92 79.55 59.74 80.47 75.43 72.57 79.73 76.72 4.81
@@ 3.3 Semantic Segmentation
The ISPRS Potsdam dataset provides both non-eroded and eroded labels, corresponding to annotations with and without object boundaries, respectively. In our experiments, we use the non-eroded labels for model evaluation and report the mean IoU for MIM- and CL-based methods. We omit LLM-based methods for semantic segmentation due to the lack of open-source MLLMs for this task. From Table 2, we can draw the following findings: 1) both MIM- and CL-based methods suffer from serious performance under image corruptions for the object detection task, with a mIoU drop of more than 10%. 2) MIM-based methods achieve much better performance than CL-based methods on clean and noisy images. This is probably because MIM-based methods can capture local details by pixel reconstruction, while CL-based methods align global visual embeddings with text features, thus losing fine-grained details. Specifically, ScaleMAE [7] achieves the best segmentation performance under image corruption, with an average mIoU of 60.02%. 3) MIM-based methods suffer from a more serious performance drop under image corruptions than CL-based methods. Specifically, GeoRSCLIP [10] achieves the best robustness across image corruptions, with a drop of 10.01% of mIoU under corruptions.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Rotate Salt Scale Translate Avg $\mathcal{R}_{\text{TP}}$
@@
SatMAE [5] ViT-L 59.51 50.18 37.39 48.23 51.15 57.89 41.83 44.95 57.52 56.02 36.07 54.81 59.09 49.59 16.67
ScaleMAE [7] ViT-L 68.92 64.37 65.43 49.96 41.84 64.86 54.89 63.89 64.49 64.51 52.12 65.45 68.38 60.02 12.91
RVSA [46] ViT-B 69.82 64.71 65.67 47.99 45.34 66.89 48.89 61.52 65.25 65.58 45.26 66.87 69.23 59.43 14.88
SatMAE++ [47] ViT-L 62.68 53.91 59.06 49.74 53.94 60.38 44.32 48.80 60.44 58.34 39.45 58.13 61.94 54.04 13.78
@@
RemoteCLIP [9] ViT-B 50.28 42.32 45.27 39.33 37.78 50.26 48.46 36.61 50.06 46.48 46.91 48.39 49.63 45.12 10.26
RemoteCLIP [9] ViT-L 56.69 54.51 52.73 43.24 51.19 50.82 45.12 50.47 51.82 53.68 38.98 54.59 56.53 50.31 11.25
GeoRSCLIP [10] ViT-B 51.44 42.89 46.41 40.37 38.64 51.35 49.79 38.56 51.15 47.89 48.24 49.28 50.87 46.29 10.01
GeoRSCLIP [10] ViT-L 56.81 54.97 52.53 43.28 41.37 50.49 42.7 49.41 51.36 53.54 36.98 54.66 56.64 48.99 13.77
@@ 3.4 Object Detection
We evaluate robustness on the corrupted images from the DIOR dataset. We report mAP for MIM- and CL-based methods. We omit LLM-based methods for the object detection task due to the lack of open-source LLMs for this task. From Table 3, we can draw the following findings: 1) both MIM- and CL-based methods suffer from serious performance under image corruptions, with a mAP drop of more than 8%. 2) MIM- and CL-based methods achieve comparable performance for object detection on clean and noisy images. RVSA [46] attains the highest mAP of 70.96% on clean images but experiences the most severe performance decline under image corruptions. 3) MIM-based and CL-based methods exhibit a similar degree of performance degradation when subjected to image corruptions. Among them, SatMAE++ [47] demonstrates the most robust detection performance under noisy conditions.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Rotate Salt Scale Translate Avg $\mathcal{R}_{\text{TP}}$
@@
SatMAE [5] ViT-L 62.30 56.84 57.86 55.80 58.36 55.38 58.44 59.34 56.92 56.60 53.76 51.58 60.90 56.82 8.81
ScaleMAE [7] ViT-L 70.20 64.80 65.98 62.50 64.46 62.58 63.82 66.10 63.08 63.44 60.50 53.08 68.26 63.22 9.94
RVSA [46] ViT-B 70.96 60.59 65.02 61.58 64.60 62.35 62.87 63.98 62.88 64.04 56.61 55.97 69.69 62.51 11.91
SatMAE++ [47] ViT-L 65.20 59.44 61.02 60.30 59.88 59.66 61.06 61.72 59.56 59.14 58.64 48.48 64.70 59.47 8.79
@@
RemoteCLIP [9] ViT-B 60.40 56.72 56.28 56.78 54.56 53.68 57.36 55.90 53.42 54.54 54.40 44.92 59.72 54.86 9.17
RemoteCLIP [9] ViT-L 70.20 66.52 66.62 63.84 65.40 63.62 63.68 66.76 62.66 63.52 59.16 57.42 68.64 63.99 8.85
GeoRSCLIP [10] ViT-B 60.20 56.28 56.04 56.08 55.46 53.38 56.92 55.50 53.38 53.98 53.48 46.98 59.32 54.73 9.09
GeoRSCLIP [10] ViT-L 69.80 66.12 65.34 65.34 64.96 63.62 62.90 66.04 62.02 62.68 56.04 57.40 68.10 63.38 9.20
@@ 3.5 Image Captioning
As shown in the upper part of Table 4, all models experience performance degradation under noisy conditions. Among them, the Falcon [54] model achieves the best overall performance, significantly outperforming other methods on both clean and corrupted images. However, it also suffers the largest performance drop of 6.28%. In contrast, RS-LLaVA [51] demonstrates the strongest robustness to image corruptions, exhibiting the smallest decrease in CLAIR score, with only a 2.03% drop. Additionally, we present results for GeoChat [12], fine-tuned on the VRSBench training set, as shown in the lower part of Table 4. The fine-tuned GeoChat model on the target dataset exhibits significantly improved performance compared to its zero-shot counterpart, with similar performance drop under corruptions.
@@
Method Backbone Clean Brightness Cloud Compression Data Gauss Gauss Haze Motion Salt Avg $\mathcal{R}_{\text{TP}}$
@@
GeoChat [12] ViT-L 41.39 40.06 40.45 37.65 40.20 39.76 38.48 40.38 39.92 37.61 39.59 4.35
SkySenseGPT [53] ViT-L 48.29 47.21 46.64 44.22 46.25 45.52 44.97 46.14 45.13 44.36 45.60 5.57
VHM [52] ViT-L 52.02 50.19 50.82 50.26 50.57 51.22 50.46 50.39 50.72 49.48 50.46 3.00
RS-LLaVA [51] ViT-L 51.30 51.15 50.43 51.78 50.54 52.01 47.84 50.57 49.88 48.12 50.26 2.03
Falcon [54] DaViT-B 61.90 59.98 60.09 57.13 59.48 57.43 56.31 59.85 59.94 51.83 58.01 6.28
GeoChatft [12] ViT-L 71.26 69.00 68.93 66.60 69.45 68.63 67.83 69.98 69.02 63.87 68.15 4.36